OpenStudy (anonymous):

factor the polynomial completely. 3x^2 - 11x - 20

OpenStudy (anonymous):

(3x +4)(x-5)

OpenStudy (anonymous):

i love factoring!! woohoo!!!

OpenStudy (anonymous):

It is important to know steps how to do it. It is shown below. 1. The first term is 3x^2. So you need something that will create 3x^2. We get (3x )(x ). 2. Now the last term is -20. So you need something that will result in -20. We need either (-5 and +4) or (-4 and +5). 3. Combine 1 and 2 and we will have 4 options (3x -5)(x +4) or (3x-4)(x+5) or (3x+5)(x-4) or (3x+4)(x-5). 4. Look at the middle term. We need to get -11x. option 1 (3x*4)+(-5*x) = 12x-5x = 7x --No. option 2 (3x*5)+(-4*x) = 15x-4x = 11x --No. option 3 (3x*-4)+(5*x) = -12x+5x = -7x --No. option 4 (3x*-5)+(4*x) = -15x+4x = -11x --Yes. 5. The answer is (3x+4)(x-5). You can double check : 3x*x =3x^2 (3x*-5)+(4*x) = -11x 4*-5 = -20 = 3x^2-11x-20 http://www.answersforquestions.com/answersforquestions/Categories/MathPolynomialFactoring/tabid/97/Default.aspx